OpenStudy (anonymous):

Write the trigonometric expression in terms of sine and cosine, and then simplify. tan θ/ sec θ − cos θ

OpenStudy (joannablackwelder):

tan(θ)=sin(θ) / cos(θ) and sec(θ) = 1/cos(θ) (I am going to use x instead of θ, it is just easier. So, sin x / cos x is the numerator so far. Lets deal with the denominator now.

OpenStudy (joannablackwelder):

1/cos x - cos x Let's combine the fraction by making like denominators. 1/cos x-cos^2 (x)/cos(x)

OpenStudy (joannablackwelder):

You with me?

OpenStudy (anonymous):

yes!

OpenStudy (joannablackwelder):

Awesome. The denominator fraction combines to (1-cos^2(x))/cosx. 1-cos^2(x) = sin^2(x)

OpenStudy (joannablackwelder):

So now, our denominator is simplified to sin^2(x)/cos(x). We can flip the denominator fraction and multiply it by our numerator and cancel, to give 1/sin(x), which is csc(x).

OpenStudy (anonymous):

thank you soooo much!!!

OpenStudy (joannablackwelder):

No problem :)